OpenStudy (anonymous):

The graph of f(x) = 2x + 1 is shown below. Explain how to find the average rate of change between x = 1 and x = 3.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@M4thM1nd

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

first find the value of function for 1 and 3 .then use this formula average rate of change=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

so that's it?

OpenStudy (anonymous):

@amirreza1870

OpenStudy (jack1):

yep, thats it

OpenStudy (anonymous):

thanks guys

OpenStudy (jack1):

if not: y = mx + c m is the gradient... which is the rate of change

OpenStudy (jack1):

in this case: y = 2x + 1 so c = 1 and m = 2...

OpenStudy (anonymous):

alright thanks

OpenStudy (jack1):

np

OpenStudy (anonymous):

The average rate of change is given by: \[a.r.c. = \frac{f(x_1)-f(x_0)}{x_1-x_0}\]

OpenStudy (anonymous):

just plug in the values